# Regret of $H_{\infty}$ Preview Controllers

### v1

@@ III BACKGROUND: $H_{\infty}$ CONTROL WITH PREVIEW
This section reviews the $p$-step $H_{\infty}$ preview controller, which provides the basis for the results in Section IV. The formulation is directly based on the standard full information $H_{\infty}$ synthesis problem described in [70, 9, 56, 18] and is included here for completeness.
III-A Full Information $H_{\infty}$ Synthesis
Consider the system (1) with full information ($p=0$). The goal of the $H_{\infty}$ problem is to design a stabilizing controller that uses measurements of the current state $x(t)$ and disturbance $d(t)$ in order to minimize the gain from $d$ to the quadratic cost (3). Since solving the exact optimal problem is difficult, it is more common to compute a $\gamma$-suboptimal $H_{\infty}$ controller. Let $\gamma>0$ be given. A controller $K$ is $\gamma$-suboptimal if it stabilizes the plant and satisfies:
$J(K,d)<\gamma^{2}\|d\|_{2}^{2}\qquad\forall d\in\ell_{2},\,d\neq 0.$ (9)
The optimal $H_{\infty}$ performance can be computed to within a desired tolerance using bisection and the solution to the $\gamma$-suboptimal $H_{\infty}$ problem. The solution to this problem is given by the next theorem.
Theorem 2.
Let $(A,B_{u},B_{d},Q,R)$ be given and satisfy conditions (i)(iv)(i)-(iv) in Theorem 1. For any $\gamma>0$, a controller $K$ that solves the $\gamma$-suboptimal $H_{\infty}$ control problem exists if and only if there exists a symmetric matrix $X$ such that:
  • 1)
[…]
  • 2)
    $H:=R+B_{u}^{\top}XB_{u}\succ 0$.
  • 3)
    $\Delta:=\gamma^{2}I-B_{d}^{\top}XB_{d}+B_{d}^{\top}XB_{u}H^{-1}B_{u}^{\top}XB_{d}\succ 0$.
Moreover, in the case that $X$ satisfies the conditions above, the optimal policy is given by:
$u_{\infty}(t)=-K_{x}x(t)-K_{d}d(t),$ (10)
where $K_{x}:=H^{-1}B_{u}^{\top}XA$ and $K_{d}:=H^{-1}B_{u}^{\top}XB_{d}$.
This result follows directly from Theorem 9.2 of [56]. Specifically, we can express the full information $H_{\infty}$ problem using the following system with a generalized error signal:
$x(t+1)$ $=Ax(t)+B_{d}d(t)+B_{u}u(t),$
$e(t)$ $=Cx(t)+Du(t),$
where $C:=\begin{bmatrix}Q^{1/2}&0\end{bmatrix}^{\top}$ and $D:=\begin{bmatrix}0&R^{1/2}\end{bmatrix}^{\top}$. The closed-loop $H_{\infty}$ norm for any controller corresponds to the induced $\ell_{2}$ gain from $d$ to $e$. Moreover, for any controller and disturbance, the cost (3) is equal to the $\ell_{2}$ norm of the error, i.e. $J(K,d)=\|e\|_{2}^{2}$. Hence, the closed-loop $H_{\infty}$ norm is less than $\gamma>0$ if and only if Equation (9) holds. Finally, Theorem 9.2 of [56] assumes $(A,B_{u},C,D)$ has no invariant zeros on the unit circle and is left invertible. Assumptions (i)-(iv) in Theorem 2 imply these assumptions hold.
III-B $H_{\infty}$ Control With Preview
Now consider the case where the controller has access to the $p$-step preview information $i_{p}(t)$ in (2). The goal of the $H_{\infty}$ preview problem is to design a stabilizing controller using this information to minimize the gain from $d$ to the quadratic cost (3). As before, the exact optimal problem is hard to solve directly, so we work with a $\gamma$-suboptimal formulation.
A suboptimal $H_{\infty}$ controller with $p$-step preview can be constructed with an augmented state that includes the future (preview) disturbances. Related preview results (both finite and infinite horizon) are given in [65, 24, 39, 29, 30, 33, 36, 34, 35, 31]. The augmented state that incorporates $p$-step future disturbances at time $t$ is:
@@ III-B $H_{\infty}$ Control With Preview
$\hat{A}_{p}$ $=\begin{bmatrix}A&B_{d}&0&\cdots&0\\ 0&0&I&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&I\\ 0&\cdots&\cdots&\cdots&0\end{bmatrix}$, $\hat{B}_{u,p}$ $=\begin{bmatrix}B_{u}\\ 0\\ \vdots\\ 0\end{bmatrix}$, $\hat{B}_{d,p}$ $=\begin{bmatrix}0\\ \vdots\\ 0\\ I\end{bmatrix}$.
@@
Thus, the preview problem reduces to a full-information $H_{\infty}$ synthesis problem on the augmented system with state $\hat{x}_{p}(t)$ and exogenous input $d(t+p)$ known at time $t$.
Assume the original state/cost matrices $(A,B_{u},B_{d},Q,R)$ satisfy the assumptions of Theorem 2. It can be shown that this implies that the augmented state/cost matrices $(\hat{A}_{p},\hat{B}_{u,p},\hat{B}_{d,p},\hat{Q}_{p},R)$ also satisfy the same assumptions for any $p\geq 0$. Hence, Theorem 2 can be applied using the augmented data. It follows that for any $\gamma>0$, a $p$-step preview controller $K$ that solves the $H_{\infty}$ suboptimal control problem (with preview) exists if and only if there exists a symmetric matrix $\hat{X}_{p}$ satisfying the conditions 1)-3) of Theorem 2 (using the augmented data). Moreover, if $\hat{X}_{p}$ satisfying the conditions exists, then the optimal $p$-preview policy is
@@
where $\hat{K}_{x,p}=H^{-1}\hat{B}_{u,p}^{\top}\hat{X}_{p}\hat{A}_{p}$ and $\hat{K}_{d,p}=H^{-1}\hat{B}_{u,p}^{\top}\hat{X}_{p}\hat{B}_{d,p}$. This policy can be further expressed directly in terms of the plant state $x(t)$ and previewed disturbances:
@@
Note that the $p$-step preview $H_{\infty}$ optimal control input, $u_{\infty,p}(t)$, can be expressed as a function of the preview information set $i_{p}(t)$. Similarly, the non-causal control input $u_{nc}(t)$ can be written in terms of the full information set $i_{nc}(t)$. This structural similarity provides useful insight into the relationship between the two controllers and helps explain the convergence results established in the next section.

### v2

@@ III BACKGROUND: $H_{\infty}$ CONTROL WITH PREVIEW
This section reviews the $p$-step $H_{\infty}$ preview controller, which provides the basis for the convergence results in Section IV. The formulation follows directly from the standard full-information $H_{\infty}$ synthesis [70, 9, 56, 18] applied to an augmented system that encodes finite disturbance preview.
A $\gamma$-suboptimal $H_{\infty}$ controller with $p$-step preview can be constructed by augmenting the state to include future disturbances. Related preview results in both finite- and infinite-horizon settings can be found in [65, 24, 39, 29, 30, 33, 36, 34, 35, 31]. The augmented state at time $t$ is defined as
@@
$\hat{A}_{p}$ $=\begin{bmatrix}A&B_{d}&0&\cdots&0\\ 0&0&I&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&I\\ 0&\cdots&\cdots&\cdots&0\end{bmatrix}$, $\hat{B}_{u,p}$ $=\begin{bmatrix}B_{u}\\ 0\\ \vdots\\ 0\end{bmatrix}$, $\hat{B}_{d,p}$ $=\begin{bmatrix}0\\ \vdots\\ 0\\ I\end{bmatrix}$. (11)
@@
Using the augmented state $\hat{x}_{p}(t)$ in (9), the $p$-step preview problem can be formulated as a full-information $H_{\infty}$ control problem on the augmented dynamics (10), where the exogenous input is $d(t+p)$ and the measured signal is $\hat{x}_{p}(t)$. Given $\gamma>0$, a controller is said to be $\gamma$-suboptimal if it stabilizes the augmented closed loop and satisfies
$J(K_{p},d)<\gamma^{2}\|d\|_{2}^{2}\qquad\forall d\in\ell_{2},\ d\neq 0.$ (13)
Theorem 2.
Fix $p\geq 0$ and let $(\hat{A}_{p},\hat{B}_{u,p},\hat{B}_{d,p},\hat{Q}_{p},R)$ be the augmented data defined in (10)–(12) and satisfy conditions (i)(iv)(i)-(iv) in Theorem 1. For any $\gamma>0$, a $\gamma$-suboptimal $p$-step preview controller exists if and only if there exists a symmetric matrix $\hat{X}_{p}$ such that:
  • 1)
[…]
  • 2)
    $H:=R+\hat{B}_{u,p}^{\top}\hat{X}_{p}\hat{B}_{u,p}\succ 0$.
  • 3)
    $\Delta_{p}\coloneq\gamma^{2}I-\hat{B}_{d,p}^{\top}\hat{X}_{p}\hat{B}_{d,p}+\hat{B}_{d,p}^{\top}\hat{X}_{p}\hat{B}_{u,p}H^{-1}\hat{B}_{u,p}^{\top}\hat{X}_{p}\hat{B}_{d,p}\succ 0$
Moreover, when these conditions hold, one $\gamma$-suboptimal preview policy is
@@
where $\hat{K}_{x,p}=H^{-1}\hat{B}_{u,p}^{\top}\hat{X}_{p}\hat{A}_{p}$ and $\hat{K}_{d,p}=H^{-1}\hat{B}_{u,p}^{\top}\hat{X}_{p}\hat{B}_{d,p}$.
Theorem 2 follows directly by applying Theorem 9.2 of [56] to the augmented plant (10) with generalized error $e(t)=\hat{C}_{p}\hat{x}_{p}(t)+Du(t)$, where $\hat{C}_{p}:=\begin{bmatrix}\hat{Q}_{p}^{1/2}&0\end{bmatrix}^{\top}$ and $D:=\begin{bmatrix}0&R^{1/2}\end{bmatrix}^{\top}$. With this choice, $J(K_{p},d)=\|e\|_{2}^{2}$ and the closed-loop induced $\ell_{2}$ gain from $d$ to $e$ is less than $\gamma$ if and only if (13) holds.
This policy can be further expressed directly in terms of the plant state $x(t)$ and previewed disturbances:
@@
Thus, $u_{\infty,p}(t)$ uses the preview information $i_{p}(t)$ in the same additive structure as the non-causal controller, but truncated to $p$ steps.